### v2

# Enabling Natural Zero-Shot Prompting on Encoder Models via Statement-Tuning

Abstract
While Large Language Models (LLMs) exhibit remarkable capabilities in zero-shot and few-shot scenarios, they often require computationally prohibitive sizes. Conversely, smaller Masked Language Models (MLMs) like BERT and RoBERTa achieve state-of-the-art results through fine-tuning but struggle with extending to few-shot and zero-shot settings due to their architectural constraints. Hence, we propose Statement-Tuning, a technique that models discriminative tasks as a set of finite statements and trains an Encoder model to discriminate between the potential statements to determine the label. We do Statement-Tuning on multiple tasks to enable cross-task generalization. Experimental results demonstrate that Statement-Tuning achieves competitive performance compared to state-of-the-art LLMs with significantly fewer parameters. Moreover, the study investigates the impact of several design choices on few-shot and zero-shot generalization, revealing that Statement-Tuning can achieve sufficient performance with modest training data and benefits from task and statement diversity for unseen task generalizability.
1 Introduction
Large Language Models (LLMs) have shown great capabilities in zero-shot and few-shot settings Radford et al. (2019); Brown et al. (2020); Artetxe et al. (2022). However, such capabilities are more difficult to observe in Encoder-only models like BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019) due to their architectural design. These models are typically pretrained in an unsupervised manner on a large corpus with a Masked Language Modeling (MLM) Devlin et al. (2019) or Discriminative Clark et al. (2020) objective and fine-tuned by adding task-specific layers to enable their usage on a particular task such as binary/multi-label classification, token/sequence classification, multiple choice, etc. These task-specific layers, thus, can not be extended effectively to new tasks in a few-shot or zero-shot manner.
In this work, we explore the feasibility of utilizing encoder models that are usually specialized for a certain task to take on various, unseen Natural Language Understanding (NLU) tasks, akin to zero-shot prompting in decoder models. One benefit of using encoder models is that they are generally more compact. Yet, encoder models have achieved state-of-the-art results on many NLU tasks through task-specific finetuning, and thus it would be interesting if zero-shot/few-shot prompting could be adapted to be compatible with encoder models to leverage their powerful NLU capabilities at more computationally feasible sizes.
To address this issue, some techniques try to reformulate various downstream tasks with a unified format resembling the pretraining objective (MLM or Discriminative pretraining), enabling few-shot transfer for encoder models Xia et al. (2022); Schick and Schütze (2021a, b); Gao et al. (2021). These approaches tend to have multiple limitations including unintuitive/difficult implementations, complex handling of multi-token labels, and being unsuitable for zero-shot generalization.
In this work, we take inspiration from multitask instruction tuning methods for decoder models Wei et al. (2022a); Sanh et al. (2022) and unified format fine-tuning methods for encoder models to propose Statement-Tuning, a novel intuitive approach for encoder-only models to generalize to zero-shot and few-shot unseen tasks using universal multitask fine-tuning. Our approach thus has the generalization ability similar to decoder models.
Figure 1: Overview of Statement-Tuning. We train an encoder to discriminate the truth value of statements from multiple tasks, then we apply it in the zero-shot setting by creating a statement for each possible target label and choosing the most likely one according to the encoder discriminator.
As seen in Figure 1, we verbalize a diverse set of NLU tasks into natural language statements, and then we fine-tune an encoder-only MLM, RoBERTa, on a universal binary sequence classification task, which we call Statement-Tuning, to assign a truth value (True or False) to any given statement. By fine-tuning encoder LLM across diverse tasks and statements, we show zero-shot generalization capabilities to unseen tasks by similarly transforming them into statements. Moreover, we show few-shot capabilities by continually fine-tuning this model with a small amount of downstream data, also formatted into statements. Statement-Tuning is capable of matching or even outperforming (32-shot and) zero-shot performance of many state-of-the-art LLMs with a fraction of the parameters.
Our ablation study shows that depending on the task, we can achieve sufficient few-shot and zero-shot generalizability with as few as 1,000 statements per training dataset or approximately 16,000 training statement examples in total, which correspond to even fewer original task examples since one example can be turned into multiple statements through different templates. Furthermore, we find that the proximity of the fine-tuning tasks to the evaluation tasks as well as the prompt and task diversity tend to have a beneficial effect on the effectiveness of Statement-Tuning.
In summary, our primary contributions are:
  1. 1.
    To the best of our knowledge we are the first to enable natural, zero-shot task generalization in encoder models by verbalizing the input into statements and fine-tuning the model to perform binary classification on the truth value of a statement.
  2. 2.
    We expose that certain emergent abilities Wei et al. (2022b) like zero-shot generalization on unseen tasks previously thought to be exclusive to decoder-based LLMs can also be observed in much smaller encoder models when we do multitask Statement-Tuning.
  3. 3.
    We explore a large number of design choices to study how Statement-Tuning benefits from the number of statement examples and statement template and task diversity in multitask Statement-Tuning.
2 Related Work
Few-shot Approaches Using Encoder-Only Models
Prompt-based fine-tuning with cloze templates or label discrimination Schick and Schütze (2021a); Gao et al. (2021) effectively utilizes encoder models for few-shot learning. Prompts provide task context beyond traditional input-output training, enabling easier adaptation. However, multi-token labels and zero-shot generalization remain challenges Schick and Schütze (2021b); Tam et al. (2021). Our work addresses this by proposing a simpler reformulation method for improved unseen task generalization via multitask training.
Verbalizers in prompts leverage label semantics for task information (e.g., Tam et al. (2021)). Similar approaches include label conditioning tasks Tam et al. (2021) and TARS Halder et al. (2020). While TARS implicitly learns label-text relations, our method uses verbalization to explicitly connect text and labels in natural language statements. This facilitates generalization to unseen tasks without further fine-tuning, as the model learns a general semantic understanding rather than a rigid input structure.
Using Masked Language Models for Sequence Scoring
In a sense, Statement-Tuning can be viewed as fine-tuning MLMs for universal NLU sequence scoring based on the truth value of a statement and the subsequent generalization to any discriminative tasks. Previous approaches have explored using MLMs for general sequence scoring for non-autoregressive generation Wang and Cho (2019); Ghazvininejad et al. (2019). Salazar et al. (2020) use MLM pseudo-log-likelihood (PLL) scores as discriminators for generated output in Automatic Speech Recognition and low-resource Machine Translation and achieve improved generation through hypothesis re-scoring using the MLM PLL scores.
Zero-shot Prompting and Multitask Tuning
LLMs excel at unseen-task/zero-shot generalization Brown et al. (2020). Building on this, recent work explores multitask training with diverse prompts for improved zero-shot performance Sanh et al. (2022); Wei et al. (2022a); Chung et al. (2022). These methods fine-tune large models on constructed datasets with various task prompts, achieving strong zero-shot results. However, effective instruction-tuned models often require billions of parameters Zhang et al. (2024), limiting their application to smaller models. Our work demonstrates achieving similar or superior generalization on smaller encoder-only models with less training data.
Figure 2: Example conversion of the MNLI task to natural language statements.
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion
Llama-2-7b-chat 7B 86.6 54.4 40.1 90.5 78.5 47.8 50.0
Mistral-7B-Instruct-v0.2 7B 89.4 73.0 41.4 88.9 82.3 57.7 55.3
Qwen1.5-7B-Chat 7B 87.0 75.5 42.1 95.3 79.7 59.1 57.8
Pythia-6.9B 6.9B 82.2 62.0 41.7 83.3 71.2 32.2 25.1
Pythia-2.8B 2.8B 79.6 68.4 41.2 77.7 69.7 12.1 35.4
Phi-2 2.7B 87.2 67.9 41.8 86.6 77.7 38.7 53.1
FlanT5-Large 770M 67.6 81.1 40.1 96.0 63.0 51.0 59.9
Qwen1.5-0.5B-Chat 500M 69.2 32.6 38.7 69.7 68.9 21.9 6.6
BART-large-mnli 406M 50.4 35.8 46.9 49.4 47.3 6.5 11.7
FlanT5-Small 60M 52.8 31.9 42.0 88.8 51.5 24.5 21.7
[…]
Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
Full/3000-shot:
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
Table 1: Comparison of our approach against many pretrained open-source Encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in Zero-shot conditions. FT stands for Full Finetuning and is included for reference. We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
3 Method: Statement-Tuning
In this section, we outline the steps involved in Statement-Tuning. First, tasks are verbalized into natural language statements. Then they are used to train the statement discriminator and derive the target label.
3.1 Task Verbalization
Any discriminative task with a finite set of targets can be verbalized into a finite set of natural language statements. Figure 2 shows the example of converting the MNLI task into statements. Similar to prompting, each task has its own statement templates, based on each possible label. The truth label for training purposes on each statement depends on whether the statement contains the correct target label or not. We outline all the statement templates used for each dataset in Appendix A.
3.2 Statement Fine-Tuning
To create the training data for statement fine-tuning, we exhaustively generate statements across 16 diverse NLP tasks using many varied statement templates per dataset: QQP Sharma et al. (2019), Winogrande Sakaguchi et al. (2020), PiQA Bisk et al. (2020), MNLI Williams et al. (2018), SNLI Bowman et al. (2015), Mintaka Sen et al. (2022), Yelp Polarity Zhang et al. (2015), WikiLingua Ladhak et al. (2020), SQuAD Rajpurkar et al. (2016), TweetEval’s Offensive task Zampieri et al. (2019), Massive FitzGerald et al. (2022); Bastianelli et al. (2020), Definite Pronoun Resolution Rahman and Ng (2012), QASC Khot et al. (2020), SciQ Johannes Welbl (2017), RACE Lai et al. (2017), and SAMSum Gliwa et al. (2019).
We explore different sample sizes for each dataset. We fine-tune RoBERTa Liu et al. (2019) with a binary sequence classification head to predict the truth value of the statements. By fine-tuning the model across diverse tasks, templates, and domains, the model should be able to generalize across unseen templates and tasks, as long as it can be phrased as a true/false statement.
3.3 Zero-Shot and Few-Shot Inference
To perform inference on statement-finetuned RoBERTa, we also need to transform the input into statements. Specifically, we exhaustively generate a statement for each possible label, as shown in Figure 1. Then, for each statement corresponding to each label, we predict the probability of such a statement being true. The final label is the statement with the highest true probability. Zero-shot inference is done by directly performing the aforementioned inference regime on the statement-finetuned RoBERTa, while K-shot inference is done by first performing continual fine-tuning on K examples of task-specific statements.
4 Experimental Setup
4.1 Evaluation
In this work, we measure our model’s task generalizability using another set of 7 diverse datasets representing a variety of unseen tasks or unseen domains during training to judge the generalizability of our approach: Balanced COPA (BCOPA) Kavumba et al. (2019); Roemmele et al. (2011), MRPC Dolan and Brockett (2005), Emotion Saravia et al. (2018), Amazon Polarity McAuley and Leskovec (2013); Zhang et al. (2015), FigQA Liu et al. (2022), StoryCloze (2016) Yang et al. (2023), and Yahoo Answers Topics Zhang et al. (2015). Among the evaluation data, BCOPA, Emotion, FigQA, StoryCloze, and Yahoo Answers Topics are unseen tasks and hence examine the cross-task generalizability; however, MRPC and Amazon Polarity represent seen tasks but in different domains and demonstrate domain generalizability.
4.2 Statement Finetuning Configurations
We statement-finetune both RoBERTa-base and RoBERTa-large across diverse NLP tasks outlined in Section 3.2. However, as statement fine-tuning expands the dataset with various templates over all possible labels, it is arguably unwise to fine-tune on all possible generated statements. Moreover, each task has a different data size, leading to unbalanced fine-tuning data. Therefore, we sample statements randomly for each task, uniformly across true and false statements. We explore sample size from 1,000 statements to 50,000 statements for each dataset. Following best practices, we encourage the invariance to phrasing in diverse statements by designing multiple statement templates per dataset (a list of all statement templates is shown in Appendix A). We also explore the effect of statement diversity during training in Section 5.2.
After statement tuning is completed, we can further continue fine-tuning the model on the target downstream dataset. Specifically, we explore various n-shot configurations: Full/3,000-shot, 1,000-shot, 500-shot, 200-shot, and 32-shot, where we use limited data from the training sets of the corresponding dataset to fine-tune our statement-tuned models. For the Full/3,000-shot case, we cap the training set at 3,000 examples, otherwise, we use the entire set (this is the case for Amazon Polarity only). For StoryCloze, there is no training set, so we just carry out 32-shot (using 32 samples from the test set for fine-tuning and evaluating on the rest) and Zero-shot experiments. As for Yahoo Answers Topic and Emotion, due to them being multi-class classification tasks, we cap the n-shot analysis at 200-shot due to the larger number of choices per example (and hence a larger number of statements per example).
4.3 Other Baselines
To assess the feasibility of our approach, we compare Statement-tuned RoBERTa base models with 125 million parameters with a range of competitive multitask fine-tuned encoder-decoder models and decoder-only LLMs spanning a parameter range from 60 million parameters to 7 billion parameters.
We include the following open-source models: Llama-2-7B-chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), QWEN1.5-7B-chat and QWEN1.5-0.5B-chat Bai et al. (2023), Pythia-6.9B and Pythia-2.8B Biderman et al. (2023), Phi-2 Li et al. (2023), FlanT5-Large and FlanT5-Small Chung et al. (2022), and BART-large-mnli Lewis et al. (2020). We use the chat/instruction-tuned version of models to allow for better prompting. We try to select models that have not seen the evaluation data to the best of our knowledge, however, the training data of many of these models is not fully outlined and there can always be the possibility of contamination. We use the standard implementation of all models in the HuggingFace transformers library Wolf et al. (2020).
We train and evaluate all the models on a configuration of 5 AMD EPYC Rome CPU cores and 2 Nvidia Tesla A100 40GB GPUs.
5 Results and Analysis
In this section, we dive deep into the results of our experimentation to derive insights about our approach.
5.1 Overall Result
Statement-Tuning Enables Effective Zero-Shot Generalization on Encoder Models
Table 1 shows zero-shot performance of statement-tuned RoBERTa. Recall that we explore various statement-tuning sizes, hence here we report the best performance across all training sizes and performance for the 4,000 and 50,000 sample sizes per dataset for the base and large models, respectively. The effect of statement tuning sample size is explored in the later part of this paper.
The result shows that the statement-tuned encoder model can achieve zero-shot generalization across unseen tasks and domains. Our approach achieves respectable performance on unseen balanced COPA, Figurative QA, StoryCloze, Yahoo Topic Answers, and Emotion. Perhaps unsurprisingly, domain generalization in sentiment analysis on Amazon Polarity and paraphrase detection on MRPC is also displayed. We also see that the larger model (RoBERTa-large) achieved much better generalization capability in general.
Comparison Against Larger Zero-Shot Models
Our approach is also competitive against other pretrained open-source encoder-decoder and decoder-only Large Language Models under zero-shot prompting. Despite having significantly fewer parameters than all the models reported (except for FlanT5-small), our approach clearly matches or exceeds many of them on the tasks reported. It is worth noting that our approach with only 125M parameters almost completely outperforms all models under or equal to 6.9B parameters on almost all tasks (except for BCOPA) and is completely dominant on FigQA and StoryCloze, both of which are unrepresented in the training data.
Furthermore, we observe generally stronger performance achieved by the Statement-Tuned RoBERTa-large models with consistently competitive performance that is on par with the 7B parameter LLMs on all the tasks. Notably, the Statement-Tuned RoBERTa-large models greatly outperform all the 7B parameter LLMs (approximately 20 times the number of parameters) on FigQA and StoryCloze, with the best performing RoBERTa-large model scoring an additional 32.6 and 10.7 points over the best performing 7B parameter LLM on the accuracy respectively.
We observe similar results in the 32-shot setting (see Appendix C). These results demonstrate the significant potential of much smaller encoder models being accurate and light alternatives to LLM zero-shot (and few-shot) prompting in natural language understanding.
Figure 3: N-shot accuracy of Statement-Tuned RoBERTa-base models across training datasets of different sizes. The x-axis denotes the number of statements per Statement-Tuning training dataset, with the number of training datasets fixed at 8. Additionally, a baseline comparison with random choice is included for reference.
5.2 Ablation Studies
Statement Average accuracy
Sample RoB-base RoB-large
1,000 62.2 69.7
2,000 64.7 72.9
3,000 66.2 73.5
4,000 67.3 73.5
5,000 66.9 73.1
10,000 66.2 73.4
20,000 66.7 73.2
40,000 66.5 73.6
50,000 58.0 74.7
Table 2: Average accuracy over all evaluation tasks when trained with different statement sample size.
Statement Finetuning Sample Size
Recall that we only perform statement fine-tuning on a sample of all possible statements from the training dataset. Here, we explore the significance of sample size per-dataset.
As shown in Table 2, a larger sample size does not always mean better overall performance for RoBERTa-base. Specifically, RoBERTa-base’s average performance plateaus around 67.3% after 4k samples and does not increase even with significantly more data. Interestingly, RoBERTa-large benefits and improves significantly from using a larger finetuning size with the best average performance observed when it is finetuned on 50,000 statements per training corpus. We hypothesize that this is due to a larger "capacity" to understand and discriminate between natural language statements which allows RoBERTa-large to benefit from more training data as opposed to RoBERTa-base which has a more limited "capacity" to develop a general semantic understanding of the truthfulness of statements. Due to computational and time constraints, we are not able to exhaust the upper limit of RoBERTa-large’s performance. Despite this, we already achieved remarkable performance with training set sizes of up to 50,000 statements. However, we leave extensive hyperparameter and training set size optimization for future work.
Figure 4: N-shot improvement of Statement-Tuned RoBERTa-base of varying training set sizes across training datasets of different sizes. The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model and the accuracy achieved by regular-finetuning of RoBERTa-base on the task. A positive Delta indicates improvement over the baseline approach.
While the statement-tuned model shows zero-shot generalization, we can further fine-tune the model with the target downstream task. As seen in Figure 3, we investigate the effect of various statement tuning training sample sizes on the base model’s accuracy across different n-shot continual fine-tuning on the 7 evaluation datasets. We observe a general trend where performance increases with more examples (shots) provided. Amazon Polarity has achieved strong performance even under a zero-shot setting, hence the improvement in a few-shot scenario is limited.
We experimented with re-purposing the model back as a standard classification model. However, performance was not as good on preliminary experimentation so we do not explore this avenue further. Detailed results can be seen in the Appendix D.
Figure 5: N-shot correlation using the average accuracy across all training set sizes and evaluation sets.
N-shot Performance
Additionally, there seems to be a general trend of improved performance by using a larger n for n-shot continual fine-tuning. However, the results seem to indicate a general trend of diminishing returns past using 200-shot finetuning. Nevertheless, it seems that a great deal of the potential performance is achieved with the zero-shot application of the approach, hence further supporting the utility of our approach when task-specific data is scarce. Additionally, as seen in Fig. 5, there appears to be a great degree of correlation between 32-shot and 0-shot model performance, indicating that observed trends in the 0-shot scenario can be informative for the few-shot case.
Effect of Statement Tuning Statement Diversity
As part of our investigation of Statement-Tuning, we would like to explore the effect of template diversity during Statement-Tuning. We hypothesize that randomly applying a larger number of different statement templates per training corpus will allow for improved performance on unseen tasks as it will help the model be more robust to the phrasing of statement templates and prevent it from forming dependencies on superficial cues in some templates.
To recall, each corpus employs several templates (see Appendix A). In this experiment, we limit each corpus to only use the maximum of N different templates, which we call Statements per Category (SPC). We Statement-Tune RoBERTa base models with a fixed training set size of 4,000 statements per corpus with a varying level of SPC.
The result can be seen in Table 3. We observe on average that increasing SPC increases performance to a certain extent before leveling off around 63.7% with an SPC of 4. The result is unsurprising, however, confirms our decision to use many statement templates per task. Detailed exploration of statement design is left to future work.
SPC BCOPA MRPC FigQA AP S-Cloze YA Topic Emotion AVG
1 72.4 65.7 59.2 92.1 86.2 33.7 46.8 62.0
2 65.0 61.2 56.9 92.3 79.4 31.9 43.8 58.4
3 69.4 70.7 59.4 91.7 85.4 37.5 48.7 63.5
4 67.6 67.1 58.6 92.5 82.1 38.5 54.8 63.7
5 70.6 71.7 58.9 91.9 79.8 40.3 46.9 63.4
Table 3: The Zero-shot performance of the base model using various degrees of SPC, where a larger SPC value indicates greater statement diversity during training. We report the average as the geometric mean of the task performance to account for the differing accuracy ranges of each task.
Statement tuning Training Evaluation
PD CR NLI QnA SA WSD IC OLI SU BCOPA MRPC FIGQA AMAZON P. StoryCloze YA Topic Emotion AVG
x x x x x x x x x 71.6 70.0 58.6 92.9 80.7 39.6 49.3 63.8
x x x x x x x x 72.4 71.0 59.2 93.1 81.2 38.3 48.5 63.8
x x x x x x x 72.2 70.7 57.0 93.1 77.7 38.0 52.5 63.6
x x x x x x 70.4 71.2 58.0 93.4 80.9 18.6 46.0 56.7
x x x x x 72.8 71.7 57.0 93.5 75.7 25.7 49.4 59.6
x x x x 72.4 71.2 54.8 60.3 74.4 28.5 53.9 57.0
x x x 70.0 67.4 59.2 78.0 75.3 36.6 40.3 58.8
Table 4: Comparison of the effect of reducing task diversity in the training of Statement-Tuning models on zero-shot accuracy on unseen datasets. The last common is the average using the geometric mean to account for the different accuracy ranges of the different evaluation sets. The total training set size remains constant at approximately 100,000 statements across all configurations.
Comparison Against Standard Fine-tuning
To observe the improvement over regular fine-tuning of RoBERTa-base, we also include Figure 4, where the y-axis, Delta, represents the improvement over regular fine-tuning for the particular n-shot. For zero-shot, we take random choice as a baseline. Generally, continually fine-tuning our model is better than fine-tuning vanilla RoBERTa under an extremely low N-shot setting. However, in some instances such as BCOPA and (to a certain extent) FigQA, we notice that even for a higher number of few-shot examples we tend to observe a benefit against regular fine-tuning.
Overall, this result demonstrates that the particular strength of our approach tends to be in the few-shot and zero-shot case, whereas, in the case of data availability, it tends to be better to directly fine-tune a RoBERTa-base model on the task. We assume that the superiority of our method in situations with limited target task data can be attributed not only to the improved generalizability gained from multitask statement tuning but also to the data augmentation effect brought by the statement tuning on few-shot target task examples, which allows the model to not only learn the relation between a text and the correct class label but also the wrong ones, thus improving data efficiency.
Effect of Statement Tuning Task Diversity
We explore how important it is to cover a variety of tasks during statement tuning. In this experiment, we group statement tuning datasets across 9 different task categories of Summarization (SU), Sentiment Analysis (SA), Question Answering (QA), Natural Language Inference (NLI), Commonsense Reasoning (CR), Paraphrase Detection (PD), Word Sense Disambiguation (WSD), Intent Classification (IC), and Offensive Language Identification (OLI). Dataset to task categories breakdown can be seen in Appendix E. We then perform statement tuning on RoBERTa-base on various subsets of tasks. To control for the total training size, we dynamically sample the data until we have 100k total statements.
Table 4 shows the 0-shot performance of the statement tuning approach when the training set size is fixed but task types are incrementally removed from the mix. We generally observe that the inclusion of more tasks seems to be overall beneficial; however, it seems to also unsurprisingly correlate with how closely the training set resembles the evaluation set. For example, we see a significant jump in accuracy on Amazon Polarity once we introduce another out-of-domain sentiment analysis data. Likewise, MRPC’s performance is already high since another paraphrase detection dataset is always introduced regardless of the configuration. Nevertheless, for truly unseen tasks, statement task diversity is beneficial, especially for multi-class classification tasks such as Yahoo Answers Topic and Emotion. In average, training task diversity seems to have an overall beneficial effect on generalizability and contributes to the ability of our approach to generalize effectively to unseen tasks.
6 Conclusion
As part of their emergent abilities, LLMs generalize to many unseen tasks and domains through few-shot and zero-shot prompting, but are prohibitively computationally expensive and difficult to adapt. To address this issue, we investigate Statement-Tuning, a novel technique for few-shot and zero-shot task generalization for encoder models. We find that this approach can match or outperform the few-shot and zero-shot prompting of many much larger decoder-only or encoder-decoder models on many tasks at a fraction of the parameters. Experimentation shows that the approach can be leveraged by training on as few as 16,000 statements. We find Statement-Tuning training task and prompt diversity to be generally helpful. We speculate that the benefits of this approach could extend beyond task generalization and could prove useful for cross-lingual task transfer, and would like to explore this in future work.
Limitations
While our approach offers advantages in computational efficiency compared to LLMs, the cost scales with the number of possible targets due to the requirement of one forward pass per label. Additionally, task-specific full fine-tuning can still achieve better performance. Furthermore, the effectiveness of our approach in generalizing to unseen tasks relies on the similarity of those tasks to the training data. Tasks with minimal relation to existing labeled datasets might see limited performance compared to tasks with higher similarity. Finally, our reliance on encoder-based models restricts its application to Natural Language Understanding tasks, excluding tasks like translation or summarization.
Ethics Statement
We affirm our commitment to more accessible and climate-aware NLP, and hope this work inspires more computationally efficient approaches to NLP. All data and models we use are publicly available. Furthermore, the success of Statement-Tuning relies on fine-tuning pretrained encoder models, which are pretrained on large datasets, and hence, Statement-Tuning is susceptible to inheriting and enforcing any harmful biases existing in the pretraining data.
Acknowledgements
Yongxin Huang is supported by HUAWEI Technologies (Ireland) Co., Ltd.
Appendix A Statement Templates
A.1 QQP Templates
Task Statement Template
QQP "{{text1}}" is a duplicate of "{{text2}}"
"{{text1}}" duplicates {{text2}}
"{{text1}}" is not a duplicate of "{{text2}}"
"{{text1}}" does not duplicate "{{text2}}"
A.2 Winogrande Templates
Task Statement Template
Winogrande In "{{sentence}}", _ is: {{option1/option2}}
Q: "{{sentence}}", A: {{option1/option2}}
The missing word in: "{{sentence}}" is {{option1/option2}}
_ in: "{{sentence}}" is {{option1/option2}}
"{{sentence}}", _ is: {{option1/option2}}
A.3 PiQA Templates
Task Statement Template
PiQA {{goal}} {{sol1/sol2}}
Goal: {{goal}}, Solution: {{sol1/sol2}}
If the goal is: {{goal}}, then the solution is: {{sol1/sol2}}
Problem: {{goal}}, Solution: {{sol1/sol2}}
A.4 MNLI and SNLI Templates
Task Statement Template
MNLI "{{text1}}" entails "{{text2}}"
{{text1}}? yes, {{text2}}
Premise: {{text1}}, Hypothesis: {{text2}}, label: Entailment
"{{text1}}" is neutral with regards to "{{text2}}"
{{text1}}? maybe, {{text2}}
Premise: {{text1}}, Hypothesis: {{text}}, label: Neutral
"{{text1}}" contradicts "{{text2}}"
{{text1}}? no, {{text2}}
Premise: {{text1}}, Hypothesis: {{text}}, label: Contradiction
A.5 Mintaka Templates
Task Statement Template
Mintaka Q: {{question}}, A: {{answerText}}
{{question}} {{answerText}}
Question: {{question}}, Answer: {{answerText}}
The answer of {{question}} is {{answerText}}
A.6 Yelp Polarity Templates
Task Statement Template
Yelp Polarity "Title: {{title}}, Content: {{content}}" has negative sentiment
{{title}} {{content}} has negative sentiment
"Title: {{title}}, Content: {{content}}", Sentiment: Negative
{{title}} {{content}} It was terrible
The sentiment in "{{title}} {{content}}" is negative
"Title: {{title}}, Content: {{content}}" has positive sentiment
{{title}} {{content}} has positive sentiment
"Title: {{title}}, Content: {{content}}", Sentiment: Positive
{{title}} {{content}} It was great
The sentiment in "{{title}} {{content}}" is positive
A.7 WikiLingua Templates
Task Statement Template
WikiLingua Passage: {{source}}, Summary: {{target}}
The summary of "{{source}}" is {{target}}
Context: {{source}}, Summary: {{target}}
Q: Summarize the following: {{source}}, A: {{target}}
The answer of "Summarize the following {{source}}" is {{target}}
A.8 SQuAD Templates
Task Statement Template
[…]
A.9 BCOPA Templates
Task Statement Template
BCOPA The cause of {{premise}} is that {{choice1/choice2}}
{{premise}} because {{choice1/choice2}}
{{premise}} due to {{choice1/choice2}}
The effect of {{premise}} is that {{choice1/choice2}}
{{premise}} therefore {{choice1/choice2}}
{{premise}}, so {{choice1/choice2}}
A.10 MRPC Templates
Task Statement Template
MRPC "{{text1}}" is a paraphrase of "{{text2}}"
[…]
{{text1}}? yes, {{text2}}
"{{text1}}" can be stated as "{{text2}}"
{{text1}}" is the same as saying "{{text2}}"
A.11 Amazon Polarity Templates
Task Statement Template
Amazon Polarity "Title: {{title}}, Content: {{content}}" has negative sentiment
{{title}} {{content}} has negative sentiment
"Title: {{title}}, Content: {{content}}", Sentiment: Negative
{{title}} {{content}} It was terrible
The sentiment in "{{title}} {{content}}" is negative
The emotions conveyed in "{{title}} {{content}}" are negative
"Title: {{title}}, Content: {{content}}" has positive sentiment
{{title}} {{content}} has positive sentiment
"Title: {{title}}, Content: {{content}}", Sentiment: Positive
{{title}} {{content}} It was great
The sentiment in "{{title}} {{content}}" is positive
The emotions conveyed in "{{title}} {{content}}" are positive
A.12 FigQA Templates
Task Statement Template
FigQA {{startphrase}} {{ending1/ending2}}
{{startphrase}} therefore {{ending1/ending2}}
startphrase: {{startphrase}}, ending: {{ending1/ending2}}
if {{startphrase}} then {{ending1/ending2}}
{{startphrase}} means {{ending1/ending2}}
A.13 StoryCloze Templates
Task Statement Template
StoryCloze {{input_sentence_1}} {{input_sentence_2}}
{{input_sentence_3}} {{input_sentence_4}}
{{sentence_quiz1/sentence_quiz2}}
A.14 Yahoo Topics Answers Templates
Task Statement Template
YA Topic {{question_title}} {{question_content}} the topic is {{topic}}
A.15 Emotion Templates
Task Statement Template
Emotion {{question_title}} {{question_content}} the topic is {{topic}}
A.16 Offensive Templates
Task Statement Template
Offensive "{{text}}". The tweet is {{label}}.
This tweet "{{text}}" is considered {{label}}.
Tweet: "{{text}}". Label: {{label}}.
"{{text}}". This text is {{label}}.
The text "{{text}}" is {{label}}.
A.17 Massive Templates
Task Statement Template
Massive The utterance "{{utt}}" is under the {{scenario}} scenario.
Utterance: "{{utt}}" Scenario: {{scenario}}
User: "{{utt}}". The best scenario for the user query is {{scenario}}.
The scenario of user’s utterance "{{utt}}" is {{scenario}}.
A.18 Definite Pronoun Resolution Templates
Task Statement Template
DPR {{sentence_with_pronoun_replaced}}
{{sentence}}. Based on the sentence, {{pronoun}} refers to {{candidates}}.
The pronoun {{pronoun}} in "{{sentence}}" is referring to {{candidates}}.
{{sentence}}. ’{{pronoun}}’ refers to {{candidates}}.
A.19 QASC Templates
Task Statement Template
QASC {{formatted_question}}. Answer: {{answer_key}}
Q: "{{formatted_question}}." A: {{answer_key}}
Question: "{{formatted_question}}." Answer: {{choices[answer_key]}}
Context: {{combined_facts}} Question: {{question}} Answer: {{choices[answer_key]}}
{{question}} Based on the passage "{{combined_facts}}", the answer if the question is "{{choices[answer_key]}}".
{{combined_facts}} {{question}} {{choices[answer_key]}}
Context: {{combined_facts}} Question: {{formatted_question}}. Answer: {{answer_key}}
{{formatted_question}}. The answer is {{answer_key}}
A.20 SciQ Templates
Task Statement Template
SciQ {{question}} {{correct_answer}}
Question: {{question}} Answer: {{correct_answer}}
{{support}} Question: {{question}} Answer: {{correct_answer}}
{{support}} According to the information, {{question}}. Answer: {{correct_answer}}.
The answer to the question {{question}}, according to "{{support}}" is {{correct_answer}}.
A.21 RACE Templates
Task Statement Template
RACE {{article}} {{question_replaced_with_answer}}
Appendix B Finetuning Setup
To finetune RoBERTa-base/RoBERTa-large on Statement-Tuning, we train for 15 epochs using an initial learning rate of 1e-06 and a weight decay of 0.01. We use a warm-up ratio of 0.1. We use 10% of the training data for validation. We use a training batch size of 16 for RoBERTa-base and a training batch size of 8 for RoBERTa-large.
Appendix C 32-shot Generalization
Table 5 shows the results of 32-shot fine-tuning on 7 target downstream datasets of our models and baselines. We observe similar trends to zero-shot setting as discussed in Section 5.1.
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze Yahoo Answers Topic Emotion
Llama-2-7b-chat 7B 91.0 67.9 42.8 95.2 82.1 61.9 54.3
Mistral-7B-Instruct-v0.2 7B 93.8 78.2 44.8 96.2 87.0 65.0 57.0
Qwen1.5-7B-Chat 7B 91.4 79.4 43.8 95.1 82.4 63.9 58.0
Pythia-6.7B 6.7B 84.6 66.9 39.2 91.6 74.0 38.3 52.0
Pythia-2.7B 2.7B 80.8 63.5 41.5 90.8 71.7 35.5 47.5
Phi-2 2.7B 90.8 74.0 44.7 93.8 81.6 58.4 58.4
FlanT5-Large 770M 66.2 78.7 39.7 75.3 59.9 38.0 34.6
Qwen1.5-0.5B-Chat 500M 73.4 56.1 38.5 84.2 68.8 36.1 31.4
BART-large-mnli 406M 52.2 32.4 42.0 50.6 51.1 7.1 10.0
FlanT5-Small 60M 52.0 32.6 41.4 75.8 50.0 9.1 9.8
[…]
Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
Full-shot:
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
Table 5: Comparison of our approach against many pretrained open-source Encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in 32-shot conditions. We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
Appendix D Regular Classification of Statement-Tuned Models
In figure 6, we visualize the relative improvement of our Statement-Tuned RoBERTa-base models regularly fine-tuned on N-shot downstream data over the regularly fine-tuned RoBERTa-base. The results are not as good as fine-tuning using statements.
Figure 6: N-shot improvement of Statement-Tuned RoBERTa-base models used for regular finetuning. The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model fine-tuned for the task directly by discarding the Statement-Tuning classification head and the accuracy achieved by regular fine-tuning of RoBERTa-base on the task. A positive Delta indicates improvement over the baseline approach.
Appendix E Task Categories Breakdown
For the statement tuning task diversity, we group corpus based on task categories as follows:
  1. 1.
    Summarization (SU): WikiLingua, SAMSum
  2. 2.
    Sentiment Analysis (SA): Yelp Polarity
  3. 3.
    Question Answering (QA): Mintaka, SQuAD, QASC, SciQ, RACE
  4. 4.
    Natural Language Inference (NLI): MNLI, SNLI
  5. 5.
    Commonsense Reasoning (CR): Winogrande, PiQA
  6. 6.
    Paraphrase Detection (PD): QQP
  7. 7.
    Word Sense Disambiguation (WSD): Definite Pronoun Resolution
  8. 8.
    Intent Classification (IC): Massive
  9. 9.
    Offensive Language Identification (OLI): Tweet Eval’s Offensive

### v3

@@
 Experimental results demonstrate that Statement-Tuning achieves competitive performance compared to state-of-the-art LLMs with significantly fewer parameters.
-Moreover, the study investigates the impact of several design choices on few-shot and zero-shot generalization, revealing that Statement-Tuning can achieve sufficient performance with modest training data and benefits from task and statement diversity for unseen task generalizability.
+Moreover, the study investigates the impact of several design choices on few-shot and zero-shot generalization, revealing that Statement-Tuning can achieve strong performance with modest training data and benefits from task and statement diversity for unseen task generalizability.
+We release all the code used to generate statement data, train and evaluate our Statement-Tuned models.
+[^1] Work In Progress
 1 Introduction
@@
 (2019) due to their architectural design.
-These models are typically pretrained in an unsupervised manner on a large corpus with a Masked Language Modeling (MLM) Devlin et al.
+These models are typically pretrained in an unsupervised manner on a large corpus with a Masked Language Modeling Devlin et al.
 (2019) or Discriminative Clark et al.
@@
 One benefit of using encoder models is that they are generally more compact.
-Yet, encoder models have achieved state-of-the-art results on many NLU tasks through task-specific finetuning, and thus it would be interesting if zero-shot/few-shot prompting could be adapted to be compatible with encoder models to leverage their powerful NLU capabilities at more computationally feasible sizes.
-To address this issue, some techniques try to reformulate various downstream tasks with a unified format resembling the pretraining objective (MLM or Discriminative pretraining), enabling few-shot transfer for encoder models Xia et al.
-(2022); Schick and Schütze (2021a, b); Gao et al.
+Yet, encoder models have achieved state-of-the-art results on many NLU tasks through task-specific fine-tuning.
+So it would be interesting if zero-shot/few-shot prompting could be adapted for encoder models to leverage their powerful NLU capabilities at more computationally feasible sizes.
+To achieve this, some techniques try to reformulate various downstream tasks with a unified format resembling the pre-training objective, enabling few-shot transfer for encoder models Schick and Schütze (2021a, b); Xia et al.
 (2022).
-These approaches tend to have multiple limitations including unintuitive/difficult implementations, complex handling of multi-token labels, and being unsuitable for zero-shot generalization.
+Without few-shot examples, the zero-shot generalization of these models relies mainly on the language modelling ability learned in the pre-training phase, not benefiting from further multitask training on diverse reformulated tasks.
 In this work, we take inspiration from multitask instruction tuning methods for decoder models Wei et al.
 (2022a); Sanh et al.
-(2022) and unified format fine-tuning methods for encoder models to propose Statement-Tuning, a novel intuitive approach for encoder-only models to generalize to zero-shot and few-shot unseen tasks using universal multitask fine-tuning.
-Our approach thus has the generalization ability similar to decoder models.
+(2022) and unified format fine-tuning methods for encoder models Xu et al.
+(2023) to propose Statement-Tuning, a novel intuitive approach for encoder-only models to generalize to zero-shot and few-shot unseen tasks through universal multitask fine-tuning with data formatted as statements.
+Our approach thus has the generalization ability similar to decoder models with a fraction of the parameters and training data.
 Figure 1: Overview of Statement-Tuning.
 We train an encoder to discriminate the truth value of statements from multiple tasks, then we apply it in the zero-shot setting by creating a statement for each possible target label and choosing the most likely one according to the encoder discriminator.
-As seen in Figure 1, we verbalize a diverse set of NLU tasks into natural language statements, and then we fine-tune an encoder-only MLM, RoBERTa, on a universal binary sequence classification task, which we call Statement-Tuning, to assign a truth value (True or False) to any given statement.
-By fine-tuning encoder LLM across diverse tasks and statements, we show zero-shot generalization capabilities to unseen tasks by similarly transforming them into statements.
+As seen in Figure 1, we verbalize a diverse set of NLU tasks into natural language statements, and then fine-tune an encoder-only MLM, RoBERTa, on a universal binary sequence classification task, which we call Statement-Tuning, to assign a truth value (True or False) to any given statement.
+By fine-tuning encoder models across diverse tasks and statements, we show zero-shot generalization capabilities to unseen tasks by similarly transforming them into statements.
 Moreover, we show few-shot capabilities by continually fine-tuning this model with a small amount of downstream data, also formatted into statements.
 Statement-Tuning is capable of matching or even outperforming (32-shot and) zero-shot performance of many state-of-the-art LLMs with a fraction of the parameters.
-Our ablation study shows that depending on the task, we can achieve sufficient few-shot and zero-shot generalizability with as few as 1,000 statements per training dataset or approximately 16,000 training statement examples in total, which correspond to even fewer original task examples since one example can be turned into multiple statements through different templates.
-Furthermore, we find that the proximity of the fine-tuning tasks to the evaluation tasks as well as the prompt and task diversity tend to have a beneficial effect on the effectiveness of Statement-Tuning.
+Our ablation study shows that depending on the task, we can achieve substantial few-shot and zero-shot generalizability with as few as 1,000 statements per training dataset or approximately 16,000 training statement examples in total, which correspond to even fewer original task examples since one example can be turned into multiple statements through different templates.
+Furthermore, we find that the statement and task diversity tend to have a beneficial effect on the performance and generalizability of Statement-Tuning.
 In summary, our primary contributions are:
   1. 1.
-    To the best of our knowledge we are the first to enable natural, zero-shot task generalization in encoder models by verbalizing the input into statements and fine-tuning the model to perform binary classification on the truth value of a statement.
+    To the best of our knowledge, we are the first to propose a combination of elaborate statement formulation and Masked Language Models as a simple and effective data/resource-efficient alternative for LLMs for zero-shot NLU task generalization.
   2. 2.
-    We expose that certain emergent abilities Wei et al.
-(2022b) like zero-shot generalization on unseen tasks previously thought to be exclusive to decoder-based LLMs can also be observed in much smaller encoder models when we do multitask Statement-Tuning.
+    We expose that the emergent ability (Wei et al., 2022b) to solve unseen tasks after training on a large volume of instruction-tuning data, which is previously thought to be exclusive to decoder-based LLMs of large-enough sizes, can also be observed in much smaller MLMs when we do multitask Statement-Tuning with a modest amount of training data.
   3. 3.
-    We explore a large number of design choices to study how Statement-Tuning benefits from the number of statement examples and statement template and task diversity in multitask Statement-Tuning.
+    We explore a large number of design choices to study how Statement-Tuning benefits from the number of statement examples and statement template and task diversity in multitask Statement-Tuning, and demonstrate the data/resource-efficiency of Statement-Tuning.
 2 Related Work
-Few-shot Approaches Using Encoder-Only Models
-Prompt-based fine-tuning with cloze templates or label discrimination Schick and Schütze (2021a); Gao et al.
-(2021) effectively utilizes encoder models for few-shot learning.
-Prompts provide task context beyond traditional input-output training, enabling easier adaptation.
-However, multi-token labels and zero-shot generalization remain challenges Schick and Schütze (2021b); Tam et al.
-(2021).
-Our work addresses this by proposing a simpler reformulation method for improved unseen task generalization via multitask training.
-Verbalizers in prompts leverage label semantics for task information (e.g., Tam et al.
-(2021)).
-Similar approaches include label conditioning tasks Tam et al.
-(2021) and TARS Halder et al.
-(2020).
-While TARS implicitly learns label-text relations, our method uses verbalization to explicitly connect text and labels in natural language statements.
-This facilitates generalization to unseen tasks without further fine-tuning, as the model learns a general semantic understanding rather than a rigid input structure.
-Using Masked Language Models for Sequence Scoring
-In a sense, Statement-Tuning can be viewed as fine-tuning MLMs for universal NLU sequence scoring based on the truth value of a statement and the subsequent generalization to any discriminative tasks.
-Previous approaches have explored using MLMs for general sequence scoring for non-autoregressive generation Wang and Cho (2019); Ghazvininejad et al.
-(2019).
-Salazar et al.
-(2020) use MLM pseudo-log-likelihood (PLL) scores as discriminators for generated output in Automatic Speech Recognition and low-resource Machine Translation and achieve improved generation through hypothesis re-scoring using the MLM PLL scores.
+Zero-Shot and Few-Shot Approaches Utilizing Label Semantics
+Various approaches have been proposed to reformulate zero-shot classification to leverage label semantics instead of indices, enabling the more generalized use of encoder models.
+In the reformulated task, textual labels are combined with the original input text, and the model should predict whether the label matches the text.
+TARS Halder et al.
+(2020) utilizes the simple concatenation of input text and label text, with no explicit connection through natural language.
+Similarly, Xu et al.
+(2023) reformulate discriminative tasks with minimal prompts, which are mostly simple concatenations of elements in the raw input.
+They find it effective for zero-shot generalization for ELECTRA-style encoder models but don’t test other encoder-only models with normal MLM training.
+Yin et al.
+(2019) propose an entailment-based approach where the input is a pair of texts: the original input text is the premise and label is converted into a hypothesis.
+However, the approach is limited by the fact that not every discriminative task can be formulated as an entailment task (in the form of a premise and a hypothesis).
+The entailment formulation also limits the training tasks to entailment datasets (MNLI, RTE, and FEVER), which often causes the model to over-rely on spurious lexical patterns, hindering generalization (Ma et al., 2021).
+Our approach Statement-Tuning serves as a more universal formulation of any discriminative task in the form of statements.
+Few-shot approaches utilizing cloze-style templates have also been proposed.
+PET Schick and Schütze (2021a) requires an ensemble of encoder models and iterative training, relies on additional unlabeled data; Improved PET variants Schick and Schütze (2021b); Tam et al.
+(2021) use complex losses to compute the probability of each token in multi-token labels.
+Our method keeps the normal sequence classification training paradigm, while in the meantime effectively leverages the label semantics, enabling straightforward zero-shot generalization on a single encoder model.
 Zero-Shot Prompting and Multitask Tuning
@@
 (2022).
-These methods fine-tune large models on constructed datasets with various task prompts, achieving strong zero-shot results.
-However, effective instruction-tuned models often require billions of parameters Zhang et al.
+These methods fine-tune large models on constructed datasets with various task prompts, achieving strong zero-shot results on unseen tasks.
+However, effective instruction-tuning often requires billions of parameters Zhang et al.
 (2024), limiting their application to smaller models.
-Our work demonstrates achieving similar or superior generalization on smaller encoder-only models with less training data.
+Ye et al.
+(2022) aim to distill this zero-shot ability in a smaller model like an LSTM through synthetic data creation using an LLM, but they create task-specific models rather than a single smaller model that is capable of generalizing.
+Our work demonstrates similar or superior generalization than LLMs using a single smaller MLM with less training data.
 Figure 2: Example conversion of the MNLI task to natural language statements.
-#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion
-Llama-2-7b-chat 7B 86.6 54.4 40.1 90.5 78.5 47.8 50.0
-Mistral-7B-Instruct-v0.2 7B 89.4 73.0 41.4 88.9 82.3 57.7 55.3
-Qwen1.5-7B-Chat 7B 87.0 75.5 42.1 95.3 79.7 59.1 57.8
-Pythia-6.9B 6.9B 82.2 62.0 41.7 83.3 71.2 32.2 25.1
-Pythia-2.8B 2.8B 79.6 68.4 41.2 77.7 69.7 12.1 35.4
-Phi-2 2.7B 87.2 67.9 41.8 86.6 77.7 38.7 53.1
-FlanT5-Large 770M 67.6 81.1 40.1 96.0 63.0 51.0 59.9
-Qwen1.5-0.5B-Chat 500M 69.2 32.6 38.7 69.7 68.9 21.9 6.6
-BART-large-mnli 406M 50.4 35.8 46.9 49.4 47.3 6.5 11.7
-FlanT5-Small 60M 52.8 31.9 42.0 88.8 51.5 24.5 21.7
-[…]
-Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
-Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
-Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
-Full/3000-shot:
-RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
-RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
-Table 1: Comparison of our approach against many pretrained open-source Encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in Zero-shot conditions.
-FT stands for Full Finetuning and is included for reference.
-We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
 3 Method: Statement-Tuning
@@
 Then they are used to train the statement discriminator and derive the target label.
-3.1 Task Verbalization
+Task Verbalization
 Any discriminative task with a finite set of targets can be verbalized into a finite set of natural language statements.
@@
 The truth label for training purposes on each statement depends on whether the statement contains the correct target label or not.
-We outline all the statement templates used for each dataset in Appendix A.
-3.2 Statement Fine-Tuning
-To create the training data for statement fine-tuning, we exhaustively generate statements across 16 diverse NLP tasks using many varied statement templates per dataset: QQP Sharma et al.
+Statement Fine-Tuning
+To create the training data for multitask statement fine-tuning, we exhaustively generate statements across 16 diverse NLP datasets (categorized into 9 tasks, see Appendix E) using many varied statement templates (see Appendix A) per dataset: QQP Sharma et al.
 (2019), Winogrande Sakaguchi et al.
@@
 (2019).
-We explore different sample sizes for each dataset.
 We fine-tune RoBERTa Liu et al.
@@
 By fine-tuning the model across diverse tasks, templates, and domains, the model should be able to generalize across unseen templates and tasks, as long as it can be phrased as a true/false statement.
-3.3 Zero-Shot and Few-Shot Inference
+Zero-Shot and Few-Shot Inference
 To perform inference on statement-finetuned RoBERTa, we also need to transform the input into statements.
-Specifically, we exhaustively generate a statement for each possible label, as shown in Figure 1.
+We randomly choose a statement template for each dataset at inference time.
+In our experiments, we show that our statement-tuning is robust to different templates.
+We exhaustively generate a statement for each possible label, as shown in Figure 1.
 Then, for each statement corresponding to each label, we predict the probability of such a statement being true.
 The final label is the statement with the highest true probability.
-Zero-shot inference is done by directly performing the aforementioned inference regime on the statement-finetuned RoBERTa, while K-shot inference is done by first performing continual fine-tuning on K examples of task-specific statements.
+Zero-shot inference is done by directly performing the aforementioned inference regime on the statement-finetuned RoBERTa, while K-shot inference is done after continual fine-tuning on K examples of task-specific statements.
+#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
+Meta-Llama-3-70B-Instruct 70B 89.0 71.3 42.0 94.7 82.7 61.9 51.8 70.5
+Llama-2-13b-chat-hf 13B 89.6 60.8 40.9 93.7 82.4 53.2 51.6 67.5
+Llama-2-7b-chat 7B 86.6 54.4 40.1 90.5 78.5 47.8 50.0 64.0
+Mistral-7B-Instruct-v0.2 7B 89.4 73.0 41.4 88.9 82.3 57.7 55.3 69.7
+Qwen1.5-7B-Chat 7B 87.0 75.5 42.1 95.3 79.7 59.1 57.8 70.9
+Pythia-6.9B 6.9B 82.2 62.0 41.7 83.3 71.2 32.2 25.1 56.8
+Pythia-2.8B 2.8B 79.6 68.4 41.2 77.7 69.7 12.1 35.4 54.9
+Phi-2 2.7B 87.2 67.9 41.8 86.6 77.7 38.7 53.1 64.7
+FlanT5-Large 770M 67.6 81.1 40.1 96.0 63.0 51.0 59.9 65.5
+Qwen1.5-0.5B-Chat 500M 69.2 32.6 38.7 69.7 68.9 21.9 6.6 43.9
+BART-large-mnli 406M 50.4 35.8 46.9 49.4 47.3 6.5 11.7 35.4
+FlanT5-Small 60M 52.8 31.9 42.0 88.8 51.5 24.5 21.7 44.7
+[…]
+RoBERTa-base (Best) 125M 75.3(0.5) 72.3(1.5) 61.4(0.6) 92.9(1.3) 79.1(1.1) 40.2(3.8) 48.5(5.1) 67.1
+RoBERTa-base (4k) 125M 72.4(0.5) 69.6(1.1) 60.7(0.9) 92.3(0.8) 78.5(2.7) 37.9(2.7) 46.6(4.3) 65.4
+RoBERTa-large (Best) 355M 85.1(0.7) 71.8(0.8) 74.2(1.4) 95.4(0.4) 92.1(0.7) 49.9(2.1) 50.7(1.4) 75.3
+RoBERTa-large (10k) 355M 85.1(0.7) 71.5(0.8) 73.0(2.4) 95.4(0.4) 91.1(0.8) 48.4(0.7) 49.1(3.2) 73.4
+Full/3000-shot:
+RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
+RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
+Table 1: Comparison of our approach against many pre-trained open-source encoder-decoder and decoder-only Large Language Models on 7 Natural Language Understanding tasks in zero-shot conditions.
+FT stands for Full Finetuning and is included for reference.
+For Statement-Tuning, we report the average across 5 training runs and 5 evaluation runs and include the average standard deviation in parenthesis.
+We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
 4 Experimental Setup
-4.1 Evaluation
-In this work, we measure our model’s task generalizability using another set of 7 diverse datasets representing a variety of unseen tasks or unseen domains during training to judge the generalizability of our approach: Balanced COPA (BCOPA) Kavumba et al.
-(2019); Roemmele et al.
-(2011), MRPC Dolan and Brockett (2005), Emotion Saravia et al.
+4.1 Evaluation Datasets
+We measure our model’s generalizability using another set of 7 diverse datasets representing a variety of unseen tasks or unseen domains: Balanced COPA (BCOPA; Kavumba et al., 2019; Roemmele et al., 2011), MRPC Dolan and Brockett (2005), Emotion Saravia et al.
 (2018), Amazon Polarity McAuley and Leskovec (2013); Zhang et al.
@@
 (2015).
-Among the evaluation data, BCOPA, Emotion, FigQA, StoryCloze, and Yahoo Answers Topics are unseen tasks and hence examine the cross-task generalizability; however, MRPC and Amazon Polarity represent seen tasks but in different domains and demonstrate domain generalizability.
+Among the evaluation data, MRPC (paraphrase identification) and Amazon Polarity (sentiment analysis) represent tasks seen during training but in different domains and demonstrate cross-domain generalizability.
+The rest are unseen tasks and hence examine the cross-task generalizability.
 4.2 Statement Finetuning Configurations
-We statement-finetune both RoBERTa-base and RoBERTa-large across diverse NLP tasks outlined in Section 3.2.
+We statement-finetune both RoBERTa-base and RoBERTa-large across diverse NLP tasks outlined in Section 3.
 However, as statement fine-tuning expands the dataset with various templates over all possible labels, it is arguably unwise to fine-tune on all possible generated statements.
@@
 Therefore, we sample statements randomly for each task, uniformly across true and false statements.
-We explore sample size from 1,000 statements to 50,000 statements for each dataset.
-Following best practices, we encourage the invariance to phrasing in diverse statements by designing multiple statement templates per dataset (a list of all statement templates is shown in Appendix A).
-We also explore the effect of statement diversity during training in Section 5.2.
-After statement tuning is completed, we can further continue fine-tuning the model on the target downstream dataset.
+In true/false statements, we also balance original classes.
+We explore sample size from 1,000 statements to 50,000 statements per dataset.
+We encourage the invariance to phrasing in diverse statements by designing multiple statement templates per dataset (a list of all statement templates is shown in Appendix A).
+Furthermore, we run the training five times to account for randomness in training data creation.
+In the evaluation, we randomly pick a template for each dataset in a single evaluation run and also repeat the evaluation five times.
+We thus report the mean and standard deviation of 5 $\times$× 5 runs to show the general task accuracy and the (in)variance to phrasing.
+We also explore the effect of statement diversity during training in Section 5.2.3.
+After multitask statement tuning is completed, we can further continue fine-tuning the model on the target downstream dataset.
 Specifically, we explore various n-shot configurations: Full/3,000-shot, 1,000-shot, 500-shot, 200-shot, and 32-shot, where we use limited data from the training sets of the corresponding dataset to fine-tune our statement-tuned models.
@@
 4.3 Other Baselines
-To assess the feasibility of our approach, we compare Statement-tuned RoBERTa base models with 125 million parameters with a range of competitive multitask fine-tuned encoder-decoder models and decoder-only LLMs spanning a parameter range from 60 million parameters to 7 billion parameters.
-We include the following open-source models: Llama-2-7B-chat Touvron et al.
+To assess the feasibility of our approach, we compare Statement-Tuned RoBERTa base models with 125 million parameters with a range of competitive multitask fine-tuned encoder-decoder models and decoder-only LLMs spanning a parameter range from 60 million parameters to 70 billion parameters.
+We include the following open-source models: Meta-Llama-3-70B-Instruct AI@Meta (2024), Llama-2-13B-chat, Llama-2-7B-chat Touvron et al.
 (2023), Mistral-7B-Instruct-v0.2 Jiang et al.
@@
 (2020).
-We use the chat/instruction-tuned version of models to allow for better prompting.
-We try to select models that have not seen the evaluation data to the best of our knowledge, however, the training data of many of these models is not fully outlined and there can always be the possibility of contamination.
-We use the standard implementation of all models in the HuggingFace transformers library Wolf et al.
-(2020).
-We train and evaluate all the models on a configuration of 5 AMD EPYC Rome CPU cores and 2 Nvidia Tesla A100 40GB GPUs.
+We use the chat/instruction-tuned version of the models to allow for better instruction following.
+We try to select models that have not seen the evaluation data to the best of our knowledge, however, the training data of many of these models is not fully outlined and there can always be the possibility of contamination Li and Flanigan (2024).
+Although these models have already been trained on a large number of instruction fine-tuning datasets, to guarantee a fair comparison as much as possible, we additionally instruction-tune a subset of the models using LoRA Hu et al.
+(2021) on the same training datasets used for Statement-Tuning.
+[^2] Due to limited computational resources, we are only able to perform this extended training and analysis on a subset of the models ranging from 500M to 13B parameters.
+Details regarding data formatting, hyper-parameters, and results are reported in Appendix H.
+We train and evaluate all the models on a configuration of 5 AMD EPYC Rome CPU cores and at most 4 Nvidia Tesla A100 40GB GPUs (we only use 4 GPUs for inference of the largest LLMs, and 1 GPU for Statement-Tuning).
+Figure 3: N-shot accuracy of Statement-Tuned RoBERTa-base models across training datasets of different sizes.
+The x-axis denotes the number of statements per Statement-Tuning training dataset, with the number of training datasets fixed.
 5 Results and Analysis
@@
 5.1 Overall Result
-Statement-Tuning Enables Effective Zero-Shot Generalization on Encoder Models
-Table 1 shows zero-shot performance of statement-tuned RoBERTa.
-Recall that we explore various statement-tuning sizes, hence here we report the best performance across all training sizes and performance for the 4,000 and 50,000 sample sizes per dataset for the base and large models, respectively.
-The effect of statement tuning sample size is explored in the later part of this paper.
-The result shows that the statement-tuned encoder model can achieve zero-shot generalization across unseen tasks and domains.
-Our approach achieves respectable performance on unseen balanced COPA, Figurative QA, StoryCloze, Yahoo Topic Answers, and Emotion.
-Perhaps unsurprisingly, domain generalization in sentiment analysis on Amazon Polarity and paraphrase detection on MRPC is also displayed.
-We also see that the larger model (RoBERTa-large) achieved much better generalization capability in general.
-Comparison Against Larger Zero-Shot Models
-Our approach is also competitive against other pretrained open-source encoder-decoder and decoder-only Large Language Models under zero-shot prompting.
-Despite having significantly fewer parameters than all the models reported (except for FlanT5-small), our approach clearly matches or exceeds many of them on the tasks reported.
-It is worth noting that our approach with only 125M parameters almost completely outperforms all models under or equal to 6.9B parameters on almost all tasks (except for BCOPA) and is completely dominant on FigQA and StoryCloze, both of which are unrepresented in the training data.
-Furthermore, we observe generally stronger performance achieved by the Statement-Tuned RoBERTa-large models with consistently competitive performance that is on par with the 7B parameter LLMs on all the tasks.
-Notably, the Statement-Tuned RoBERTa-large models greatly outperform all the 7B parameter LLMs (approximately 20 times the number of parameters) on FigQA and StoryCloze, with the best performing RoBERTa-large model scoring an additional 32.6 and 10.7 points over the best performing 7B parameter LLM on the accuracy respectively.
-We observe similar results in the 32-shot setting (see Appendix C).
-These results demonstrate the significant potential of much smaller encoder models being accurate and light alternatives to LLM zero-shot (and few-shot) prompting in natural language understanding.
-Figure 3: N-shot accuracy of Statement-Tuned RoBERTa-base models across training datasets of different sizes.
-The x-axis denotes the number of statements per Statement-Tuning training dataset, with the number of training datasets fixed at 8.
-Additionally, a baseline comparison with random choice is included for reference.
+Table 1 shows zero-shot performance of statement-tuned RoBERTa and baselines.
+Recall that we explore various statement-tuning sizes, hence here we report the best performance across all training sizes and performance for the 4,000 and 10,000 sample sizes per dataset for the base and large models, respectively.
+The effect of statement tuning sample size is explored in Section 5.2.1.
+Statement-Tuning Enables Effective Zero-Shot Generalization on Masked Language Models.
+The result shows that the multitask statement-tuned encoder model can achieve zero-shot generalization across unseen tasks and domains.
+On BCOPA (unseen task) and Amazon Polarity (unseen domain), our zero-shot statement-tuned models even achieve accuracies on par with models fine-tuned on the full datasets.
+We also see that the larger model (RoBERTa-large) achieved much better generalization than the base model in general.
+Comparison Against Larger Zero-Shot Models.
+Our approach is also competitive against other pre-trained open-source encoder-decoder and decoder-only Large Language Models under zero-shot prompting.
+Despite having significantly fewer parameters than all the models reported (except for FlanT5-small), our approach matches or exceeds many of them on the tasks reported.
+In average, our best Statement-Tuned RoBERTa-large models with only 355M parameters outperform the best performing LLM (Qwen1.5-7B-Chat) by 4.4 and the largest LLM (Meta-Llama-3-70B-Instruct, with approximately 200 times the number of parameters) by 4.8.
+It is worth noting that our RoBERTa-base models with only 125M parameters almost completely outperforms all models under or equal to 6.9B parameters (except for FlanT5-Large) on all tasks (except for BCOPA).
+Our models are dominant on FigQA and StoryCloze, both of which are unrepresented in the training data, with the best performing RoBERTa-large model scoring an additional 32.2 and 9.4 points over Llama3-70B-Instruct on the accuracy respectively.
+We observe similar results in the 32-shot setting (see Appendix C) and when the LLMs are additionally instruction-tuned on the same data (see Appendix H).
+These results demonstrate the capabilities of much smaller encoder models as being accurate and light (in terms of parameters; for speed comparison see Appendix G) alternatives to LLM zero-shot (and few-shot) prompting in natural language understanding.
 5.2 Ablation Studies
-Statement Average accuracy
-Sample RoB-base RoB-large
-1,000 62.2 69.7
-2,000 64.7 72.9
-3,000 66.2 73.5
-4,000 67.3 73.5
-5,000 66.9 73.1
-10,000 66.2 73.4
-20,000 66.7 73.2
-40,000 66.5 73.6
-50,000 58.0 74.7
-Table 2: Average accuracy over all evaluation tasks when trained with different statement sample size.
-Statement Finetuning Sample Size
-Recall that we only perform statement fine-tuning on a sample of all possible statements from the training dataset. Here, we explore the significance of sample size per-dataset.
-As shown in Table 2, a larger sample size does not always mean better overall performance for RoBERTa-base.
-Specifically, RoBERTa-base’s average performance plateaus around 67.3% after 4k samples and does not increase even with significantly more data.
-Interestingly, RoBERTa-large benefits and improves significantly from using a larger finetuning size with the best average performance observed when it is finetuned on 50,000 statements per training corpus.
-We hypothesize that this is due to a larger "capacity" to understand and discriminate between natural language statements which allows RoBERTa-large to benefit from more training data as opposed to RoBERTa-base which has a more limited "capacity" to develop a general semantic understanding of the truthfulness of statements.
-Due to computational and time constraints, we are not able to exhaust the upper limit of RoBERTa-large’s performance.
-Despite this, we already achieved remarkable performance with training set sizes of up to 50,000 statements.
-However, we leave extensive hyperparameter and training set size optimization for future work.
-Figure 4: N-shot improvement of Statement-Tuned RoBERTa-base of varying training set sizes across training datasets of different sizes.
-The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model and the accuracy achieved by regular-finetuning of RoBERTa-base on the task.
+Statement Sample Average accuracy
+RoB-base RoB-large
+1,000 63.0 71.1
+2,000 62.7 71.0
+3,000 63.1 73.3
+4,000 65.4 72.9
+5,000 64.7 72.2
+10,000 64.3 73.4
+20,000 64.9 68.6
+40,000 64.1 72.0
+50,000 58.5 68.2
+Table 2: Average accuracy over all evaluation tasks when trained with different statement sample size per dataset.
+Figure 4: N-shot improvement of Statement-Tuned RoBERTa-base of varying training set sizes over standard fine-tuning.
+The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model and the accuracy achieved by regular fine-tuning of RoBERTa-base on the task.
 A positive Delta indicates improvement over the baseline approach.
-While the statement-tuned model shows zero-shot generalization, we can further fine-tune the model with the target downstream task.
-As seen in Figure 3, we investigate the effect of various statement tuning training sample sizes on the base model’s accuracy across different n-shot continual fine-tuning on the 7 evaluation datasets.
-We observe a general trend where performance increases with more examples (shots) provided.
-Amazon Polarity has achieved strong performance even under a zero-shot setting, hence the improvement in a few-shot scenario is limited.
-We experimented with re-purposing the model back as a standard classification model.
-However, performance was not as good on preliminary experimentation so we do not explore this avenue further.
-Detailed results can be seen in the Appendix D.
-Figure 5: N-shot correlation using the average accuracy across all training set sizes and evaluation sets.
-N-shot Performance
-Additionally, there seems to be a general trend of improved performance by using a larger n for n-shot continual fine-tuning.
+5.2.1 Statement Finetuning Sample Size
+Recall that we only perform statement fine-tuning on a sample of all possible statements from the training dataset.
+Here, we explore the effect of sample size per dataset in the multitask statement fine-tuning on both zero-shot and few-shot performance.
+Zero-Shot.
+As shown in Table 2, with only 1k samples per datasets, we can already reach 96% of the best performance, which is obtained with 4k samples on RoBERTa-base and 10k samples on RoBERTa-large, showing the sample-efficiency of statement-tuning.
+For RoBERTa-base, introducing more data after 4,000 samples does not further improve the accuracy on downstream tasks and even causes a decrease.
+RoBERTa-large benefits from a larger fine-tuning data size with the best average performance observed when the statement number per training set increases to 10,000, more than doubling the optimal sample size of RoBERTA-base.
+We hypothesize that this is due to a larger capacity to understand and discriminate between natural language statements which allows RoBERTa-large to benefit from more training data as opposed to RoBERTa-base, which has a more limited capacity to develop a general semantic understanding of the truthfulness of statements.
+Nonetheless, we also observe a decrease in RoBERTa-large’s performance after 10k samples.
+The existence of a point of diminishing returns in both models when it comes to Statement-Tuning training data sizes indicates that too many samples may lead to overfitting to the training tasks, which affects the generalizability.
+This is consistent with previous observation that excessive instruction tuning can hurt the performance of smaller models (8B and below) on held-out tasks due to more limited model capacity Wei et al.
+(2022a).
+Few-Shot.
+While the statement-tuned model shows zero-shot generalization, we can further fine-tune the model on the target downstream task.
+As seen in Figure 3, we investigate the effect of both the multitask statement-tuning sample size and the number of shots from the target tasks on the n-shot performance on the 7 evaluation datasets.
+When increasing the multitask statement-tuning sample size, we observe a trend in n-shot performance similar to the general zero-shot performance shown in Table 2.
+For example, the optimal data size is achieved at around 4k~5k on COPA, Emotion and Yahoo, and there is an apparent drop in accuracy across different shot numbers and tasks when increasing the sample size from 40k to 50k.
+There turns out to be a high degree of correlation among all n-shot and 0-shot performance (see Figure 6), indicating that observed trends in the 0-shot scenario can be informative for the few-shot cases.
 However, the results seem to indicate a general trend of diminishing returns past using 200-shot finetuning.
 Nevertheless, it seems that a great deal of the potential performance is achieved with the zero-shot application of the approach, hence further supporting the utility of our approach when task-specific data is scarce.
-Additionally, as seen in Fig.
-5, there appears to be a great degree of correlation between 32-shot and 0-shot model performance, indicating that observed trends in the 0-shot scenario can be informative for the few-shot case.
-Effect of Statement Tuning Statement Diversity
-As part of our investigation of Statement-Tuning, we would like to explore the effect of template diversity during Statement-Tuning.
-We hypothesize that randomly applying a larger number of different statement templates per training corpus will allow for improved performance on unseen tasks as it will help the model be more robust to the phrasing of statement templates and prevent it from forming dependencies on superficial cues in some templates.
-To recall, each corpus employs several templates (see Appendix A).
-In this experiment, we limit each corpus to only use the maximum of N different templates, which we call Statements per Category (SPC).
-We Statement-Tune RoBERTa base models with a fixed training set size of 4,000 statements per corpus with a varying level of SPC.
-The result can be seen in Table 3.
-We observe on average that increasing SPC increases performance to a certain extent before leveling off around 63.7% with an SPC of 4.
-The result is unsurprising, however, confirms our decision to use many statement templates per task.
-Detailed exploration of statement design is left to future work.
+5.2.2 Comparison with Standard Fine-Tuning
+To observe the improvement over regular fine-tuning of RoBERTa-base, we also include Figure 4, where the y-axis, Delta, represents the improvement over regular fine-tuning for the particular n-shot.
+For zero-shot, we take random choice as the baseline.
+Generally, continually fine-tuning our model is better than fine-tuning vanilla RoBERTa under an extremely low N-shot setting.
+However, in some instances such as BCOPA and (to a certain extent) FigQA, we notice that even for a higher number of few-shot examples, we tend to observe a benefit against regular fine-tuning.
+Our approach is recommended in extreme few-shot and zero-shot scenarios.
+When more data is available, directly fine-tuning RoBERTa-base is better.
+Our method’s good performance with limited data can be attributed to improved generalizability from multitask statement tuning and data augmentation effect of statements generated from few-shot examples, which enhances data efficiency.
 SPC BCOPA MRPC FigQA AP S-Cloze YA Topic Emotion AVG
-1 72.4 65.7 59.2 92.1 86.2 33.7 46.8 62.0
-2 65.0 61.2 56.9 92.3 79.4 31.9 43.8 58.4
-3 69.4 70.7 59.4 91.7 85.4 37.5 48.7 63.5
-4 67.6 67.1 58.6 92.5 82.1 38.5 54.8 63.7
-5 70.6 71.7 58.9 91.9 79.8 40.3 46.9 63.4
+1 91.0(1.7) 74.8(2.2) 49.8(4.8) 59.1(0.5) 58.4(17.1) 78.0(3.4) 41.2(2.1) 62.6(6.9)
+[…]
+[…]
+[…]
+[…]
 Table 3: The Zero-shot performance of the base model using various degrees of SPC, where a larger SPC value indicates greater statement diversity during training.
 We report the average as the geometric mean of the task performance to account for the differing accuracy ranges of each task.
+Each value is a mean over 5 evaluation runs and we include the standard deviation in the parentheses.
 Statement tuning Training Evaluation
 PD CR NLI QnA SA WSD IC OLI SU BCOPA MRPC FIGQA AMAZON P. StoryCloze YA Topic Emotion AVG
-x x x x x x x x x 71.6 70.0 58.6 92.9 80.7 39.6 49.3 63.8
-x x x x x x x x 72.4 71.0 59.2 93.1 81.2 38.3 48.5 63.8
-x x x x x x x 72.2 70.7 57.0 93.1 77.7 38.0 52.5 63.6
-x x x x x x 70.4 71.2 58.0 93.4 80.9 18.6 46.0 56.7
-x x x x x 72.8 71.7 57.0 93.5 75.7 25.7 49.4 59.6
-x x x x 72.4 71.2 54.8 60.3 74.4 28.5 53.9 57.0
-x x x 70.0 67.4 59.2 78.0 75.3 36.6 40.3 58.8
+x x x x x x x x x 71.0(0.9) 65.7(3.1) 59.8(1.0) 90.7(1.3) 75.1(3.8) 36.8(2.7) 46.2(3.8) 61.2(2.7)
+x x x x x x x x 69.8(2.6) 65.9(6.1) 60.4(0.4) 91.2(0.7) 79.8(1.7) 29.4(3.7) 47.1(0.4) 60.0(3.0)
+x x x x x x x 70.0(0.3) 64.2(7.9) 59.3(0.2) 92.0(0.3) 70.5(6.4) 31.2(2.3) 49.4(3.0) 59.6(4.1)
+x x x x x x 68.7(2.1) 64.6(6.8) 58.8(0.7) 91.3(0.5) 77.3(7.0) 18.7(3.0) 53.4(4.9) 56.5(4.4)
+x x x x x 70.2(0.5) 67.0(4.6) 59.4(0.8) 91.8(0.1) 73.4(6.7) 20.5(3.4) 52.2(3.0) 57.2(3.5)
+x x x x 70.0(1.2) 67.4(2.5) 59.2(0.3) 78.0(12.0) 75.3(10.1) 36.6(3.3) 40.3(2.2) 58.8(6.2)
+x x x 50.2(1.7) 40.6(8.5) 50.9(1.7) 55.9(5.5) 50.2(6.6) 3.8(1.8) 7.2(2.8) 26.0(4.8)
 Table 4: Comparison of the effect of reducing task diversity in the training of Statement-Tuning models on zero-shot accuracy on unseen datasets.
-The last common is the average using the geometric mean to account for the different accuracy ranges of the different evaluation sets.
+The last column is the average using the geometric mean to account for the different accuracy ranges of the different evaluation sets.
 The total training set size remains constant at approximately 100,000 statements across all configurations.
-Comparison Against Standard Fine-tuning
-To observe the improvement over regular fine-tuning of RoBERTa-base, we also include Figure 4, where the y-axis, Delta, represents the improvement over regular fine-tuning for the particular n-shot.
-For zero-shot, we take random choice as a baseline.
-Generally, continually fine-tuning our model is better than fine-tuning vanilla RoBERTa under an extremely low N-shot setting.
-However, in some instances such as BCOPA and (to a certain extent) FigQA, we notice that even for a higher number of few-shot examples we tend to observe a benefit against regular fine-tuning.
-Overall, this result demonstrates that the particular strength of our approach tends to be in the few-shot and zero-shot case, whereas, in the case of data availability, it tends to be better to directly fine-tune a RoBERTa-base model on the task.
-We assume that the superiority of our method in situations with limited target task data can be attributed not only to the improved generalizability gained from multitask statement tuning but also to the data augmentation effect brought by the statement tuning on few-shot target task examples, which allows the model to not only learn the relation between a text and the correct class label but also the wrong ones, thus improving data efficiency.
-Effect of Statement Tuning Task Diversity
-We explore how important it is to cover a variety of tasks during statement tuning.
-In this experiment, we group statement tuning datasets across 9 different task categories of Summarization (SU), Sentiment Analysis (SA), Question Answering (QA), Natural Language Inference (NLI), Commonsense Reasoning (CR), Paraphrase Detection (PD), Word Sense Disambiguation (WSD), Intent Classification (IC), and Offensive Language Identification (OLI).
-Dataset to task categories breakdown can be seen in Appendix E.
-We then perform statement tuning on RoBERTa-base on various subsets of tasks.
-To control for the total training size, we dynamically sample the data until we have 100k total statements.
-Table 4 shows the 0-shot performance of the statement tuning approach when the training set size is fixed but task types are incrementally removed from the mix.
-We generally observe that the inclusion of more tasks seems to be overall beneficial; however, it seems to also unsurprisingly correlate with how closely the training set resembles the evaluation set.
-For example, we see a significant jump in accuracy on Amazon Polarity once we introduce another out-of-domain sentiment analysis data.
-Likewise, MRPC’s performance is already high since another paraphrase detection dataset is always introduced regardless of the configuration.
-Nevertheless, for truly unseen tasks, statement task diversity is beneficial, especially for multi-class classification tasks such as Yahoo Answers Topic and Emotion.
-In average, training task diversity seems to have an overall beneficial effect on generalizability and contributes to the ability of our approach to generalize effectively to unseen tasks.
+5.2.3 Effect of Statement Diversity
+As part of our investigation of Statement-Tuning, we would like to explore the effect of template diversity during Statement-Tuning.
+We hypothesize that randomly applying a larger number of different statement templates per training corpus will allow for improved performance on unseen tasks, as it will make the model more robust to the phrasing of statement templates and prevent it from relying on superficial cues in certain templates.
+In our main experiments, each dataset employs several templates (see Appendix A).
+In this experiment, we limit each corpus to only use the maximum of N different templates, which we call Statements per Category (SPC).
+We statement-tune RoBERTa base models with a fixed training set size of 4,000 statements per training corpus with a varying level of SPC.
+Table 3 shows that though BCOPA and StoryCloze benefits from a larger SPC, increasing SPC doesn’t always boost average task performance, with the highest being 62.6 at SPC 1.
+However, average standard deviation drops significantly from 6.9% to 3.6% when SPC increases from 1 to 2, reaching a low of 3.0% at SPC 3.
+This suggests that more template diversity improves stability and consistency.
+Therefore, using at least 2 different templates is recommended for better robustness.
+5.2.4 Effect of Task Diversity
+We examine the importance of task variety in Statement-Tuning.
+Our Statement-Tuning datasets can be grouped into 9 task categories: Summarization (SU), Sentiment Analysis (SA), Question Answering (QA), Natural Language Inference (NLI), Commonsense Reasoning (CR), Paraphrase Detection (PD), Word Sense Disambiguation (WSD), Intent Classification (IC), and Offensive Language Identification (OLI).
+See Appendix E for the dataset breakdown.
+We perform statement tuning on RoBERTa-base with various task subsets, dynamically sampling data to maintain 100k total statements.
+Table 4 shows the zero-shot performance of the statement tuning approach with a fixed training set size but varying task types.
+Average performance increases from 26.0 to 61.2 as the number of tasks is increased from the minimum of 3 to the maximum of 9.
+Robustness also improves, shown by a decrease in average standard deviation from 4.8% to 2.7%.
+This demonstrates that increasing training task diversity can boost performance and reduce variance.
+Unsurprisingly, the inclusion of the Sentiment Analysis task substantially improves the performance on Amazon Polarity from the same task category.
+Another related task, Emotion, also shows a large increase after adding the SA task.
+Although the inclusion of SA and WSD hurts the performance on a dissimilar task, Yahoo Answer Topic, the accuracy is recovered after adding the more related Intent Classification task, and reaches the highest 36.8 when training with all tasks.
+However, the enhancement of downstream tasks does not always come from similar training tasks.
+More interestingly, adding the QA task leads to a significant jump in the performance of all evaluation tasks.
+Though Paraphrase Identification is always included in the training, MRPC still benefits from the QA task, reflected by a great improvement of 26.8.
+Both related and unrelated training tasks can have a positive effect on the downstream tasks, highlighting the value of task diversity in multitask statement-tuning.
+Sometimes adding an unrelated task causes a performance drop on certain datasets, e.g.
+StoryCloze after adding IC, but including more tasks alleviates the problem, again confirming the advantage of task diversity.
 6 Conclusion
-As part of their emergent abilities, LLMs generalize to many unseen tasks and domains through few-shot and zero-shot prompting, but are prohibitively computationally expensive and difficult to adapt.
+As part of their emergent abilities, LLMs generalize to many unseen tasks/domains through few-shot and zero-shot prompting, but are prohibitively computationally expensive and difficult to adapt.
 To address this issue, we investigate Statement-Tuning, a novel technique for few-shot and zero-shot task generalization for encoder models.
-We find that this approach can match or outperform the few-shot and zero-shot prompting of many much larger decoder-only or encoder-decoder models on many tasks at a fraction of the parameters.
+We find that this approach can match/outperform the few-shot and zero-shot prompting of many much larger decoder-only or encoder-decoder models on many tasks at a fraction of the parameters.
 Experimentation shows that the approach can be leveraged by training on as few as 16,000 statements.
-We find Statement-Tuning training task and prompt diversity to be generally helpful.
+We find training task and statement template diversity to be generally helpful.
 We speculate that the benefits of this approach could extend beyond task generalization and could prove useful for cross-lingual task transfer, and would like to explore this in future work.
@@
 While our approach offers advantages in computational efficiency compared to LLMs, the cost scales with the number of possible targets due to the requirement of one forward pass per label.
-Additionally, task-specific full fine-tuning can still achieve better performance.
-Furthermore, the effectiveness of our approach in generalizing to unseen tasks relies on the similarity of those tasks to the training data.
-Tasks with minimal relation to existing labeled datasets might see limited performance compared to tasks with higher similarity.
-Finally, our reliance on encoder-based models restricts its application to Natural Language Understanding tasks, excluding tasks like translation or summarization.
+That being said, it is still possible to apply our method in extreme multi-class classification because we do not have to use all possible statements with all possible labels for training, as our model is trained exactly to generalize to unseen labels by learning the relation of label semantics and the input text.
+Additionally, task-specific full fine-tuning can still achieve better performance in the presence of more training data.
+Therefore, we recommend the use of our approach in the low-resource/no-resource scenario.
+Furthermore, our method can be sensitive to the Statement-Tuning training set size and other hyperparameters, hence some exploration of ideal hyperparameters may be required before employing Statement-Tuned models.
+In addition, we limit our analysis only to English, it would be interesting to observe whether the technique enables cross-lingual transfer but we leave this to future work.
+Finally, our reliance on encoder-based models restricts its application to Natural Language Understanding tasks, excluding tasks like translation or abstractive summarization.
 Ethics Statement
@@
 Table 5 shows the results of 32-shot fine-tuning on 7 target downstream datasets of our models and baselines. We observe similar trends to zero-shot setting as discussed in Section 5.1.
-#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze Yahoo Answers Topic Emotion
-Llama-2-7b-chat 7B 91.0 67.9 42.8 95.2 82.1 61.9 54.3
-Mistral-7B-Instruct-v0.2 7B 93.8 78.2 44.8 96.2 87.0 65.0 57.0
-Qwen1.5-7B-Chat 7B 91.4 79.4 43.8 95.1 82.4 63.9 58.0
-Pythia-6.7B 6.7B 84.6 66.9 39.2 91.6 74.0 38.3 52.0
-Pythia-2.7B 2.7B 80.8 63.5 41.5 90.8 71.7 35.5 47.5
-Phi-2 2.7B 90.8 74.0 44.7 93.8 81.6 58.4 58.4
-FlanT5-Large 770M 66.2 78.7 39.7 75.3 59.9 38.0 34.6
-Qwen1.5-0.5B-Chat 500M 73.4 56.1 38.5 84.2 68.8 36.1 31.4
-BART-large-mnli 406M 52.2 32.4 42.0 50.6 51.1 7.1 10.0
-FlanT5-Small 60M 52.0 32.6 41.4 75.8 50.0 9.1 9.8
+#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze Yahoo Answers Topic Emotion Avg
+Meta-Llama-3-70B-Instruct 70B 95.2 78.9 46.1 96.6 86.5 66.4 58.8 75.5
+Llama-2-13b-chat-hf 13B 93.2 71.6 44.3 95.7 84.9 61.1 56.6 72.5
+Llama-2-7b-chat 7B 91.0 67.9 42.8 95.2 82.1 61.9 54.3 70.7
+Mistral-7B-Instruct-v0.2 7B 93.8 78.2 44.8 96.2 87.0 65.0 57.0 74.6
+Qwen1.5-7B-Chat 7B 91.4 79.4 43.8 95.1 82.4 63.9 58.0 73.4
+Pythia-6.7B 6.7B 84.6 66.9 39.2 91.6 74.0 38.3 52.0 63.8
+Pythia-2.7B 2.7B 80.8 63.5 41.5 90.8 71.7 35.5 47.5 61.6
+Phi-2 2.7B 90.8 74.0 44.7 93.8 81.6 58.4 58.4 71.7
+FlanT5-Large 770M 66.2 78.7 39.7 75.3 59.9 38.0 34.6 56.1
+Qwen1.5-0.5B-Chat 500M 73.4 56.1 38.5 84.2 68.8 36.1 31.4 55.5
+BART-large-mnli 406M 52.2 32.4 42.0 50.6 51.1 7.1 10.0 35.0
+FlanT5-Small 60M 52.0 32.6 41.4 75.8 50.0 9.1 9.8 38.7
 […]
-Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
-Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
-Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
+Our Approach: RoBERTa-base (4k) 125M 75.0(0.6) 70.0(1.9) 60.3(1.0) 92.4(0.8) 79.7(3.5) 38.2(2.5) 45.4(3.2) 65.9
+Our Approach: RoBERTa-large (Best) 355M 85.1(0.8) 71.5(1.9) 74.7(1.8) 95.3(0.8) 91.9(0.2) 50.2(2.2) 49.8(1.4) 74.1
+Our Approach: RoBERTa-large (10k) 355M 85.1(0.8) 71.5(0.8) 72.6(1.8) 95.3(0.3) 91.0(1.0) 48.2(0.8) 48.4(3.6) 73.1
 Full-shot:
-RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
-RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
+RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
+RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
 Table 5: Comparison of our approach against many pretrained open-source encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in 32-shot conditions.
-We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
+We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
 Appendix D Regular Classification of Statement-Tuned Models
@@
 Appendix E Task Categories Breakdown
-For the statement tuning task diversity, we group corpus based on task categories as follows:
+For the statement tuning task diversity, we group datasets based on task categories as follows (evaluation datasets are underlined):
   1. 1.
@@
   2. 2.
-    Sentiment Analysis (SA): Yelp Polarity
+    Sentiment Analysis (SA): Yelp Polarity, Amazon Polarity
   3. 3.
@@
   6. 6.
-    Paraphrase Detection (PD): QQP
+    Paraphrase Detection (PD): QQP, MRPC
   7. 7.
@@
     Offensive Language Identification (OLI): Tweet Eval’s Offensive
+  10. 10.
+    Sentence Completion: BCOPA, StoryCloze
+  11. 11.
+    Emotion Recognition: Emotion
+  12. 12.
+    Topic Classification: Yahoo Answer Topic
+  13. 13.
+    Nonliteral Reasoning: FigQA
+Appendix F N-shot Correlation
+Figure 6: N-shot correlation using the average accuracy across all training set sizes and evaluation sets.
+Figure 6 shows the correlation of accuracies achieved in different N-shot settings with various shot numbers.
+Appendix G Inference Speed Comparison
+We report the average examples/sec processed for each of the datasets in Table 6.
+It is important to note that all models are run on a single GPU, except for Meta-Llama-3-70B-Instruct and Llama-2-13B-chat which were run on 4 and 2 GPUs, respectively.
+Model BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
+Qwen1.5-0.5B-Chat 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
+phi-2 1.4 1.4 1.4 1.4 1.5 1.4 1.4 1.4
+Meta-Llama-3-70B-Instruct* 2.9 1.3 3.2 0.9 4.9 1.1 2.1 2.3
+flan-t5-large 8.2 13.2 13.2 13.2 13.2 13.2 13.2 12.5
+Llama-2-13b-chat-hf* 8.7 5.7 12.8 4.3 15.7 4.4 6.9 8.3
+Our Approach (roberta-large) 9.3 14.5 15.0 15.0 14.7 3.1 5.1 11.0
+bart-large-mnli 9.7 14.1 14.0 14.2 14.1 13.7 13.8 13.4
+pythia-6.9b 12.0 0.6 4.6 0.4 0.6 2.2 0.4 3.0
+Llama-2-7b-chat-hf 12.5 0.6 4.6 0.4 0.6 2.3 0.5 3.1
+Mistral-7B-Instruct-v0.2 12.8 0.5 2.7 0.3 0.5 1.7 0.4 2.7
+pythia-2.8b 13.6 16.7 24.9 15.2 27.2 15.1 20.9 19.1
+flan-t5-small 13.9 39.2 39.1 39.3 39.4 39.3 39.3 35.6
+Our Approach (roberta-base) 17.9 49.8 50.0 49.8 49.9 10.3 17.0 34.9
+Table 6: The average examples per second processed by each model on each task. * indicates that the model required the use of more than one GPU.
+Llama-2-13b Qwen1.5-7B Pythia-6.9B Pythia-2.9B Phi-2 Qwen1.5-0.5B
+#Parameters 13B 7B 6.9B 2.9B 2.7B 500M
+Quantization 8bit 4bit 4bit 4bit 4bit 4bit
+Sequence Length 4096 2048 2048 2048 2048 2048
+lora_r 32 32 32 32 32 32
+lora_alpha 16 64 64 64 64 64
+lora_dropout 0.05 0.05 0.05 0.05 0.05 0.05
+adam_beta2 0.999 0.999 0.95 0.95 0.95 0.999
+adam_epsilon 1e-8 1e-8 0.00001 0.00001 0.00001 1e-8
+max_grad_norm none none 1.0 1.0 1.0 none
+optimizer adamw_bnb_8bit adamw_torch adamw_torch adamw_torch adamw_torch adamw_torch
+gradient acc. 4 4 4 4 4 4
+micro batch size 2 1 1 1 1 1
+lr_scheduler cosine cosine cosine cosine cosine cosine
+learning_rate 0.0002 0.0002 0.00001 0.00001 0.000003 0.0002
+Table 7: Instruction Tuning Hyperparameters
+Appendix H Statement-Tuning Data Instruction-Tuned Decoders
+As seen in Table 8, we report the comparison of our approach with several decoders that were instruct-tuned using an instruction-tuning dataset created using the same training corpora used for Statement-Tuning.
+Training details are outlined in Table 7.
+Furthermore, the templates used to form instructions are based on those used for Flan Wei et al.
+(2022a) (we make the dataset available here: Anonymized Repository).
+The models tested are a subset of the ones reported in Table 1 due to the time and computational expense of instruction-tuning and hardware limitations.
+However, even when the LLM models are instruction-tuned on the same data as the Statement-Tuned RoBERTa models, we observe similar trends where performance of the RoBERTa-base model tends to match the performance of all models up to 6.9B parameters on all tasks except for BCOPA.
+Furthermore, the RoBERTa-large model approaches or exceeds performance on all tasks for models with 7B+ parameters.
+The same trend of dominating performance on FigQA and StoryCloze is observed.
+#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
+Llama-2-13b 13B 89.6 60.8 40.9 93.7 82.4 53.2 51.6 67.5
+Qwen1.5-7B 7B 87.2 78.9 41.4 94.8 75.7 47.8 56.5 68.9
+Pythia-6.9B 6.9B 82.8 68.1 40.0 71.7 71.5 16.4 27.5 54.0
+Pythia-2.9B 2.9B 79.6 67.9 40.3 77.2 69.7 21.2 30.6 55.2
+Phi-2 2.7B 87.2 68.1 41.7 85.6 77.8 38.4 53.5 64.6
+Qwen1.5-0.5B 500M 72.4 68.4 39.4 49.8 67.6 33.2 72.4 57.6
+[…]
+Our Approach: RoBERTa-base (4k) 125M 72.4(0.5) 69.6(1.1) 60.7(0.9) 92.3(0.8) 78.5(2.7) 37.9(2.7) 46.6(4.3) 65.4
+Our Approach: RoBERTa-large (Best) 355M 85.1(0.7) 71.8(0.8) 74.2(1.4) 95.4(0.4) 92.1(0.7) 49.9(2.1) 50.7(1.4) 75.3
+Our Approach: RoBERTa-large (10k) 355M 85.1(0.7) 71.5(0.8) 73.0(2.4) 95.4(0.4) 91.1(0.8) 48.4(0.7) 49.1(3.2) 73.4
+Full/3000-shot:
+RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
+RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
+Table 8: Comparison of our approach against many pretrained open-source encoder-Decoder and Decoder-only Instruction-tuned Pretrained Large Language Models on 7 Natural Language Understanding tasks in Zero-shot conditions.
+FT stands for Full Finetuning and is included for reference.
+For Statement-Tuning, we report the average across 5 training runs and 5 evaluation runs and include the average standard deviation in parenthesis.
+We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.